OpenStudy (anonymous):

factor 4x^2-4 completely

OpenStudy (anonymous):

(2x)^2-2^2

myininaya (myininaya):

@aJ01 wrote your expression like that to hint to you what you need to factor.

OpenStudy (anonymous):

that still does not make sense to me

OpenStudy (e.mccormick):

Ever heard about the difference of square?

OpenStudy (anonymous):

hint: if a = 2x and b = 2 (a + b)(a - b) = a^2 + b^2

OpenStudy (anonymous):

4(x^2-1) 4(x^2-1^2) since 1^2=1 4(x-1)(x+1)